OpenStudy (anonymous):

Can someone please explain this inverse, 1) f(x) = x^4 - 3x^2 + 18 -I know there's no inverse but why and how do i prove this?-

OpenStudy (anonymous):

I think there is an inverse, mainly because quartics are solvable.

OpenStudy (anonymous):

There is? how would i even start...im at a lose with this problem.

OpenStudy (anonymous):

Ok, we have y = x^4 + 3x^2+18 inverting, we have x = y^4+3y^2+18 solve for y for the inverse

OpenStudy (anonymous):

ah...idk why i kept tying factor it... but thank you =)

OpenStudy (anonymous):

it doesnt factor, it doesnt solve, its discriminate is negative